OpenStudy (anonymous):

8. Find the sum of the infinite geometric sequence. 1 -1/5 + 1/25 - 1/125 . . .

OpenStudy (paxpolaris):

\[\Large \sum_{n=0}^{\infty}\left( -\frac 15 \right)^n\]

OpenStudy (anonymous):

= 5/4?

OpenStudy (paxpolaris):

no, less then 1 .... 5/6

OpenStudy (paxpolaris):

if it converges... \[\Large\sum_{n=0}^{\infty}ar^n = {a \over 1-r}\]